A rectangle's width is six times its length. The perimeter of the rectangle is two hundred thirty-eight feet. What is the area of the rectangle?
Answer by vleith(2517)

Area = L*W
W = 6L
Perimeter = 2L + 2W = 238
P = 2L + 2(6L)
P = 14L
238 = 14L
17=L
W = 6L = 102
A=WL = 102*17 = 1734